OpenStudy (anonymous):

Tell me if I am right Find the vertex,line of symmetry,and the minumum/maximum value of the quadratic function f(x)=-2x^2+2x+4 x-corrdinate is 9/2 y is 1/2 equation of line of symmetry is x=1/2 maximum is 9/2

OpenStudy (anonymous):

i think you have the x and y backwards

OpenStudy (anonymous):

so the line of symmetry would be x=9/2?

OpenStudy (anonymous):

no i confused you i think the axis of symmetry is what you wrote \[x=\frac{1}{2}\] and the maximum you also have right. it is \[\frac{9}{2}\]

OpenStudy (anonymous):

those two are correct

OpenStudy (anonymous):

oh, ok. i just have the x and y switched around then?

OpenStudy (anonymous):

but you wrote "the x - coordinate is 9/2" but that is the y - coordinate

OpenStudy (anonymous):

first one is x, second is y. so vertex is \[(\frac{1}{2},\frac{9}{2})\]

OpenStudy (anonymous):

good job though!